### v2

# Anant-Net: Breaking the Curse of Dimensionality with Scalable and Interpretable Neural Surrogate for High-Dimensional PDEs

Abstract
High-dimensional partial differential equations (PDEs) arise in diverse scientific and engineering applications but remain computationally intractable due to the curse of dimensionality. Traditional numerical methods struggle with the exponential growth in computational complexity, particularly on hypercubic domains, where the number of required collocation points increases rapidly with dimensionality. Here, we introduce Anant-Net, an efficient neural surrogate that overcomes this challenge, enabling the solution of PDEs in high dimensions. Unlike hyperspheres, where the internal volume diminishes as dimensionality increases, hypercubes retain or expand their volume (for unit or larger length), making high-dimensional computations significantly more demanding. Anant-Net efficiently incorporates high-dimensional boundary conditions and minimizes the PDE residual at high-dimensional collocation points. To enhance interpretability, we integrate Kolmogorov–Arnold networks into the Anant-Net architecture. We benchmark Anant-Net’s performance on several linear and nonlinear high-dimensional equations, including the Poisson, Sine-Gordon, and Allen-Cahn equations, demonstrating high accuracy and robustness across randomly sampled test points from high-dimensional space. Importantly, Anant-Net achieves these results with remarkable efficiency, solving 300-dimensional problems on a single GPU within a few hours. We also compare Anant-Net’s results for accuracy and runtime with other state-of-the-art methods. Our findings establish Anant-Net as an accurate, interpretable, and scalable framework for efficiently solving high-dimensional PDEs.
Keywords: High-dimensional PDEs; Scalable and Interpretable Architecture; Physics-Informed Neural Surrogates;
Kolmogorov–Arnold networks (KAN).
1 Introduction
Physics-informed deep learning (PIDL) represents a rapidly advancing framework that integrates known governing physical laws, typically formulated as PDEs, into the training process of deep neural networks. In contrast to conventional data-driven models that rely solely on observational data, PIDL incorporates physical constraints to guide learning, thereby enhancing generalization, reducing data dependence, and improving interpretability. This synthesis of physics and deep learning has demonstrated broad applicability in solving forward and inverse problems across scientific and engineering domains, particularly in scenarios involving limited, noisy, or deceptive data. Key methodologies under the PIDL umbrella include physics-informed neural networks (PINNs) [1, 2, 3, 4], which embed PDE constraints via automatic differentiation; sparse identification of nonlinear dynamics (SINDy) [5, 6], which infers governing equations by promoting sparsity in learned representations; and neural operators [7, 8, 9, 10, 11], which approximate solution operators across function spaces to model families of PDEs. These approaches are particularly well-suited for high-dimensional problems, where traditional numerical solvers suffer from the curse of dimensionality.
High-dimensional PDEs are integral to various scientific and engineering domains, including quantum mechanics, financial mathematics, and optimal control. Their solutions provide crucial insights into complex, multi-scale phenomena that cannot be accurately captured using lower-dimensional approximations. However, solving these equations efficiently remains a significant challenge due to the curse of dimensionality, the exponential growth in computational complexity and data requirements as the number of dimensions increases. This issue is particularly pronounced in numerical approximations, making high-dimensional PDEs computationally intractable with conventional methods. In this work, we introduce Anant-Net (with Anant meaning infinite), a novel neural architecture designed to mitigate the curse of dimensionality and enable the efficient solution of high-dimensional PDEs critical to real-world applications. Our approach provides an efficient framework for handling large datasets, whose size typically scales exponentially with dimensionality. While purely data-driven models might seem a viable alternative, they often struggle with excessive data requirements, making them impractical for high-dimensional problems. To address this, we leverage physics-informed neural surrogates, incorporating the underlying PDE as an additional constraint. This integration significantly reduces the dependency on large datasets, offering a more scalable and computationally efficient framework for solving high-dimensional PDEs.
Figure 1: Variation of the volume of a hypersphere/hypercube in high-dimensional spaces, highlighting the motivation for Anant-Net in addressing high-dimensional PDEs in hypercube domain. Here, $R$ denotes the radius of the hypersphere, and $L$ represents the edge length of the hypercube.
In high-dimensional spaces, the volume of a hypersphere decreases as dimensionality increases when the radius $R>1,<1,=1$, eventually approaching zero. Conversely, the volume of a hypercube grows exponentially with dimension if the edge length $L>1$, while it remains constant at $L=1$; see figure 1. Since the volume of a hypersphere shrinks with increasing dimensions, enforcing data constraints exactly within this region makes the problem more easier. In contrast, for a hypercube with $L>1$, the exponentially growing volume poses significant computational challenges, making it difficult to efficiently solve high-dimensional PDEs.
Range 21 Dims. 99 Dims. 300 Dims.
$[0,1]$ ($L=1$) 0.4 $\pm$ 2.17e-03 0.7 $\pm$ 2.67e-03 0.5 $\pm$ 3.44e-03
$[-1,1]$ ($L>1$) 1.49 $\pm$ 1.49e-02 7.09 $\pm$ 3.71e-02 6.00 $\pm$ 3.81e-02
Table 1: Anant-Net: Mean and standard deviation (computed over 10 random realizations) of the % relative $L_{2}$ testing error on randomly selected test points for the high-dimensional Poisson equation.
Table 1 presents the mean and standard deviation of percentage relative $L_{2}$ error (testing error) of Anant-Net evaluated on randomly selected test points for the high-dimensional Poisson equation. These testing errors show that Anant-Net can accurately solve high-dimensional PDEs.
1.1 Related Work
In the context of high-dimensional PDEs, two primary factors determine the efficacy of any method based on scientific machine learning: (1) the sampling strategy for data or collocation points, and (2) the computation of gradients for residual losses via automatic differentiation, which are used as soft constraints in training physics-informed models. It is evident that a significant number of existing methods represent a trade-off between these factors, which ultimately influences their scalability and suitability for solving real-world problems of practical relevance. While various methods exist for solving high-dimensional PDEs, this work focuses on neural surrogates that not only leverage available data but also incorporate physical laws, ensuring the consistency of the solutions with underlying governing principles. In recent years, the challenge of solving high-dimensional PDEs has been addressed by several researchers through the use of neural surrogate methods. Han et al. [12] introduced a deep neural network framework that reformulates high-dimensional PDEs as backward stochastic differential equations. However, their approach relied on approximating the gradient of the solution using a neural network, which introduced scalability issues, particularly in scenarios involving long-time integration. This limitation was subsequently addressed by Raissi et al. [13], who proposed an alternative method that directly models the solution of the PDE with a neural network, rather than its gradient. The efficacy of this improvement was demonstrated through its application to the Black-Scholes-Barenblatt and Hamilton-Jacobi-Bellman equations in 100 dimensions. Hu et al. [14] introduced a stochastic dimensions gradient descent (SDGD) framework based on physics-informed neural networks to solve high-dimensional PDEs on hypercubes. This method stochastically samples the dimensions of the PDE and computes the residuals only in the selected sample dimensions, rather than performing the computation across all dimensions. Their approach significantly reduced the computational burden on hardware, leading to notable speed-ups and improved accuracy compared to other contemporary methods for high-dimensional PDEs. However, their method bypassed the need for data or boundary point sampling by directly enforcing boundary conditions within the model, particularly for hypersphere, where the internal volume approaches zero as the number of dimensions increases (see Figure 1). In the similar line of work, Hu et al.[15] proposed Monte Carlo fractional PINN (MC-fPINN) and its extension, MC-tfPINN for tempered fractional PINN (an extension to fPINN [16]) to solve extremely high-dimensional problems. Further, an alternate approach was proposed by the same authors [17] extended fractional PINN to solve high-dimensional Fokker-Planck-Lévy (FPL) equations in high-dimensions. Adaptive training methods have also been proposed as supplementary strategies to enhance the accuracy of physics-informed neural network models for high-dimensional PDEs [18].
Wang et al. [19] introduced tensor neural networks for solving high-dimensional PDEs, wherein the independent input features are distributed across sub-networks to construct a low-rank basis representation of the input space. The solution is then mapped as a superposition of these low-dimensional basis vectors. The advantages of their method were demonstrated through experiments for both homogeneous and non-homogeneous boundary value problems up to 20 dimensions. Despite its superior accuracy, the method faces scalability limitations as the number of sub-networks increases with dimension, leading to significant computational burdens on hardware. More recently, De et al. [20] applied physics-informed extreme learning machines to high-dimensional PDEs, framing the training of random neural networks as a least-squares problem to determine the unknown weights in the final layer, while keeping the weights in the previous hidden layers fixed. Their approach achieved both superior accuracy and low computational cost for high-dimensional problems, with training demonstrated for up to 100 dimensions. Unlike automatic differentiation, their method used finite differences to compute gradients for high-dimensional PDEs. Among many strategies that were proposed to tackle the curse of dimensionality, one of the most promising strategy is to avoid the discretizations in space/time or using a grid-based format for high-dimensional data. To this end, Darbon et al. [21] proposed two frameworks based on deep neural networks to approximate the viscosity solutions of the Hamiltonian Jacobi PDEs under certain reasonable assumptions on the PDE. Richter [22] et al. proposed a tensor train format for parabolic PDEs wherein the in-built orthogonality of their proposed format enabled fast and efficient optimization in stochastic backward formulations of the PDE. Meng et al. [23] proposed an architecture that fuses SympNet with PINNs to tackle curse of dimensionality in constrained optimal multi-agent path planning problems. They investigated various formulations of the loss function to arrive at the one with least violation and later post-processed with pseudo-spectral method to achieve spectral accuracy for their path trajectories. In a similar line of work, Varghese et al.[24] SympNet was extended to Symplectic Graph Neural Networks (SympGNNs) that can effectively handle system identification in high-dimensional Hamiltonian systems such as a 2000-particle molecular dynamics simulation in a two-dimensional Lennard-Jones potential.
Several studies have leveraged the weak formulation of PDEs to train deep learning models, particularly for high-dimensional problems. A pioneering approach in this direction is the Deep Ritz method [25], which utilizes the variational formulation of PDEs by minimizing an energy functional. This method, designed to address high-dimensional PDEs, employs a ResNet-based architecture and is trained using stochastic gradient descent. It systematically samples data from both the interior and the boundary of the domain, distinguishing itself from approaches such as that of Hu et al. [14], which circumvent boundary data sampling. Building on this framework, the Weak Adversarial Network (WAN) [26] was introduced to solve high-dimensional PDEs by reformulating the problem of finding a weak solution as an operator norm minimization task. In contrast to conventional numerical methods such as finite difference schemes, WAN demonstrated computational efficiency, stability, and the ability to mitigate the curse of dimensionality. The methodology employs generative adversarial network (GAN) framework, where the weak solution and test functions are parameterized by deep neural networks, termed the primal and adversarial networks, respectively. More recently, the WAN framework has been extended through the introduction of XNODE-WAN [27], which replaces the primal network with a Neural Ordinary Differential Equation (Neural ODE) model. This enhancement has resulted in improved computational speed and accuracy, particularly for time-dependent high-dimensional PDEs. These approaches systematically sample data across both the interior and the boundary of the computational domain to enhance the robustness of their solutions.
1.2 Our Contributions
The main contributions of this work are as follows:
  • We propose the Anant-Net architecture, a scalable and efficient neural architecture for solving high-dimensional PDEs defined on hypercubic domains. The network is designed to maintain high accuracy while addressing the computational challenges inherent to high-dimensional settings, particularly for hypercubes with side length $L>1$. The Anant-Net framework leverages tensor product structures to compute PDE solutions across the specified batch-size dimensions and performs dimension-wise sweeps over the full batch to handle high-dimensional problems. This tensor product formulation enables the efficient processing of large volumes of training data and collocation points, which is important for scaling to high dimensions.
  • We identify data sampling as an important and resource-intensive aspect when solving high-dimensional PDEs. To address this, we propose a sampling-based strategy that reveals the relationship between accuracy, speedup, and the volume of boundary/data points as the PDE dimension increases. Moreover, we develop an efficient framework to handle large datasets, enabling scalable training and evaluation.
  • To mitigate the high cost associated with computing derivatives in high-dimensional spaces, Anant-Net avoids exhaustive differentiation. Instead, it strategically samples a subset of dimensions and applies automatic differentiation selectively, thereby preserving the complex structure of the governing PDE while significantly reducing computational complexity.
  • We evaluate the performance of Anant-Net by reporting test accuracy on randomly sampled points from high-dimensional space after training. Specifically, we present the mean and variance of the test errors. Anant-Net demonstrates the ability to solve PDEs in up to 300 dimensions within a few hours on a single GPU, highlighting its scalability and computational efficiency.
  • We also introduce an interpretable variant of Anant-Net based on the Kolmogorov–Arnold Network, called as Anant-KAN, enabling further insights into the learned solution structures in high-dimensional settings.
  • We benchmark Anant-Net against several state-of-the-art methods, reporting comparative results in terms of runtime as well as the mean and variance of relative testing error (%), and observe that Anant-Net consistently outperforms the alternatives.
The paper is organized as follows: Section 2 presents a detailed description of the Anant-Net architecture, including the methodologies employed for structuring and integrating training data and collocation points. Section 3 provides numerical results for a representative set of high-dimensional boundary value problems, along with a comparative analysis against existing state-of-the-art architectures. Finally, Section 4 highlights the key findings of this work with a detailed discussion.
2 Methodology: Anant-Net Architecture
Separable physics-informed neural networks (SPINNs) [28] were recently proposed as an architectural framework for solving high-dimensional PDEs. In this approach, the independent features of an $d$-dimensional PDE are distributed across $n$-body subnetworks. Leveraging the structure of the architecture, SPINNs employ Einstein summation to efficiently combine the outputs of the $n$ independent subnetworks to construct the overall solution field. This strategy has been shown to perform exceptionally well in terms of both accuracy and computational efficiency for moderate-dimensional problems. However, the method exhibits limitations in scalability, as its performance degrades for problems with dimensionality greater than five. This is primarily due to the computational burden associated with naively applying Einstein summation over a increasing number of subnetworks, which becomes intractable as dimensions increases. Despite this limitation, a key strength of the SPINN framework lies in its ability to process a large volume of training data and collocation points, achieving high throughput without incurring additional computational or memory overhead, even in the absence of specialized hardware.
Figure 2: Schematic representation of general Anant-Net architecture with batch-size $B$, highlighting its core components and computational flow for solving high-dimensional PDEs.
The primary objective of the proposed architecture is to enable the scalable solution of high-dimensional PDEs. As discussed in the previous section, several prominent network architectures that report strong performance on high-dimensional PDE benchmarks often circumvent this challenge by embedding boundary conditions directly into the training process, thereby avoiding the explicit sampling of interior and boundary data points. In contrast, the present work addresses this limitation by introducing the Anant-Net architecture capable of efficiently processing large volumes of training and collocation data, a necessary condition for robust performance at scale. Our method not only accommodates the data-intensive demands inherent to high-dimensional problems but also consistently achieves high accuracy on randomly chosen test data for different problems.
Figure 2 shows the general architecture of Anant-Net. The architecture comprises a number of body networks equal to the batch size. Let $u\in\mathbb{R}^{d}$ denote the solution to a $d$-dimensional PDE, representing the unknown quantity to be approximated by Anant-Net. For brevity, we consider a batch size of three, the Anant-Net prediction of the solution field, denoted by $\hat{u}_{\theta}$ , is expressed as:
$\hat{u}_{\theta}(X_{1},X_{2},X_{3})=f_{1}^{(\theta_{1})}(X_{1})\otimes f_{2}^{(\theta_{2})}(X_{2})\otimes f_{3}^{(\theta_{3})}(X_{3}),$ (1)
where the inputs are partitioned as $X_{1}=\{x_{1},\dots,x_{d/3}\}$, $X_{2}=\{x_{d/3+1},\dots,x_{2d/3}\}$ , and $X_{3}=\{x_{2d/3+1},\dots,x_{d}\}$, such that $X_{1}\cap X_{2}\cap X_{3}=\emptyset$. The indices $p\in\mathrm{D}_{1}=\{1,2,\ldots,d/3\},q\in\mathrm{D}_{2}=\{d/3+1,2,\ldots,2d/3\},r\in\mathrm{D}_{3}=\{2d/3+1,2,\ldots,d\}$ , corresponding to the number of batch elements, are stochastically sampled from the set of active dimensions, i.e., $\{x_{p},x_{q},x_{r}\}\subset X=X_{1}\cup X_{2}\cup X_{3}$, which sweeps over all dimensional indices. The network parameters are given by $\theta=\bigcup\theta_{i},\ \forall i$ (all body networks), where $f_{j}^{(\theta_{i})}$ denotes the output vector of the $i^{\text{th}}$ body network. Although there is no strict ordering required for partitioning the dimension set $\mathrm{D}=\mathrm{D}_{1}\cup\mathrm{D}_{2}\cup\mathrm{D}_{3}$, we adopt an ascending-order partitioning strategy for all experiments presented in this paper for the sake of consistency. The number of partitioned subsets is arbitrary and can be increased by proportionally increasing the number of body networks in the Anant-Net architecture. However, for brevity, we fix the total number of partitioned sets; and correspondingly, the number of body networks, to three in all experiments discussed herein. Algorithm 1 gives the details of training data and collocation points for the Anant-Net architecture. It is also important to note that modifying the number of body networks by changing the number of partitioned dimension subsets must be reflected in the data loader, as described in Algorithm 1.
1: Consider a $d$-dimensional PDE where $\mathrm{D}=\{1,2,3,\dots d\}$ is the set of all the dimensional indices
2: Partitioning the dimensions into $B=\{p,q,r\}$ sets (equal to the batch size), corresponding to the $B=3$ body networks.
3: $\{p,q,r\}$ are the stochastically sampled active dimensions (the dimension across which the solution has a non-zero gradient) and $\mathrm{D}$\$\{p,q,r\}$ are inactive dimensions.
repeat
       Sample $\{p,q,r\}$ randomly
      Yield $\mathcal{G}_{\{p,q,r\}}^{\text{data}},\leavevmode\nobreak\ \mathcal{G}_{\{p,q,r\}}^{\text{coll.}},\leavevmode\nobreak\ u_{\{p,q,r\}}^{\text{data}},\leavevmode\nobreak\ \forall\hskip 5.69046ptp,q,r\in\mathrm{D}$
      (Here $\mathcal{G}$ represent the grid)
until Total number of batches ($d/B$) are covered;
Algorithm 1 Training Data and Collocation points loader for Anant-Net architecture.
For a general PDE of the form as shown below,
$\mathcal{L}u=f\text{ in }\Omega^{d}\text{ and }\mathcal{B}u=g\text{ on }\partial\Omega^{d}$ (2)
where, $u\in\mathbb{R}^{d}$, $\mathcal{L}$ and $\mathcal{B}$ are operators acting on the interior, $\Omega$ and boundary, $\partial\Omega$ respectively. $\mathcal{B}u=g$ is the boundary constraint. The above equation can be recast as an equivalent optimization problem with the following loss function:
$\mathcal{J}(\theta)=\lambda_{r}\|\mathcal{L}\hat{u}_{\theta}-f\|_{2}^{2}+\lambda_{b}\|\mathcal{B}\hat{u}_{\theta}-g\|_{2}^{2}$ (3)
where, $\lambda_{r}$ and $\lambda_{b}$ are the weights on PDE residual loss and boundary loss respectively, which are suitably chosen to avoid any gradient imbalance between the two objectives while training the model for high dimensional PDEs; see [29] and [30] for more details. $\hat{u}_{\theta}$ is the output of the Anant-Net that approximate the solution of a PDE. Algorithm 2 shows the details of Anant-Net training procedure. Appendix C provides a detailed description of all the symbols used in this work.
1: In general, for $B$-batch size, the prediction of the solution field ($\hat{u}_{\theta}$) via Anant-Net architecture is given by:
$\hat{u}_{\theta}(X_{1},X_{2},\ldots,X_{B})=f_{1}^{(\theta_{1})}(X_{1})\otimes f_{2}^{(\theta_{2})}(X_{2})\otimes\leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \otimes f_{B}^{(\theta_{B})}(X_{B}).$
For $B=3$, the active dimension indices $p,q,r$, corresponding to the batch size, are stochastically sampled such that all dimensions are swept without repetition within a single epoch.
repeat
       2: Sample $\mathcal{G}_{\{p,q,r\}}^{\text{data}}$, $\mathcal{G}_{\{p,q,r\}}^{\text{coll.}}$, and the corresponding dimension set by using Algorithm 1
      3: Compute data loss ($\mathcal{J}_{\text{data}}$) :
$\mathcal{J}_{\text{data}}\left(\mathcal{G}_{\{p,q,r\}}^{\text{data}},u^{\text{data}};\theta\right)=\left\|\hat{u}_{\theta}\left(\mathcal{G}_{\{p,q,r\}}^{\text{data}}\right)-u^{\text{data}}\left(\mathcal{G}_{\{p,q,r\}}^{\text{data}}\right)\right\|_{2}^{2}.$
      4: Compute PDE residual loss ($\mathcal{J}_{\text{Res}}$) on the sampled (active) dimensions $\{p,q,r\}$ within a hypercube:
$\mathcal{J}_{\text{Res}}\left(\mathcal{G}_{\{p,q,r\}}^{\text{coll.}};\theta\right)=\biggl{\|}\sum_{k\in\{p,q,r\}}\mathcal{L}_{k}\hat{u}_{\theta}(x_{1},x_{2},\dots x_{d})-f(x_{1},x_{2},\dots x_{d})\biggr{\|}_{2}^{2}.$
Note that, since the PDE loss is computed only over the active dimensions, all inactive dimensions are assigned constant values based on their respective spatial locations in a computational domain.
      5: Compute total loss ($\mathcal{J}_{\text{Total}}$):
$\mathcal{J}_{\text{Total}}\left(\mathcal{G}_{\{p,q,r\}}^{\text{data}},\mathcal{G}_{\{p,q,r\}}^{\text{coll.}};\theta\right)=\lambda_{b}\cdot\mathcal{J}_{\text{data}}\left(\mathcal{G}_{\{p,q,r\}}^{\text{data}},u^{\text{data}};\theta\right)+\lambda_{r}\cdot\mathcal{J}_{\text{Res}}\left(\mathcal{G}_{\{p,q,r\}}^{\text{coll.}};\theta\right).$
      6: Compute gradient of the total loss, update the parameters based on the average gradient using the suitable optimizer.
until i $\leq$ Total number of batches ($d/B$);
Algorithm 2 Anant-Net training procedure
2.1 Anant-Net’s Data Structures: Balancing Training Efficiency and Testing Flexibility
For the Anant-Net architecture, it is important to highlight the non-trivial differences between the training and testing procedures. During training, the data structure follows a separable (or tensor-array) format, as described in Algorithm 1. In this format, only $B$ selected dimensions are active, meaning they exhibit active gradients, while the remaining $(d-B)$ dimensions remain fixed in an arbitrary $d$-dimensional problem. This tensor-based format offers enhanced memory efficiency, which helps prevent memory bottlenecks even in problems with extremely large dimensionality. Furthermore, the tensor-based training framework enables the data to be handled as hypercubes, fully leveraging the Anant-Net architecture described in Algorithm 2. However, using a tensor-array structure during the testing phase would render the model overly rigid, thereby limiting its applicability to more practical scenarios involving randomly selected dimensions. To address this, the testing phase employs a distinct approach, wherein the data structure adopts a vector-array format, consisting of randomly sampled testing points (and not a grid). This format allows unrestricted selection of dimensions from an arbitrary $d$-dimensional space during inference.
2.2 Theoretical Analysis of Computational Complexity
Here, we consider space complexity as the function of the number of trainable parameters in a given model architecture. For a fully-connected neural network of depth, $D$ (layers) and width, $W$ (neurons per layer), the space complexity of the model can be approximated as $\mathcal{O}(W^{2}D)$ assuming no bias terms. However, for a separable neural network (such as SPINN[28]) with $B$ body networks, the space complexity is given by $\mathcal{O}(B\tilde{W}^{2}\tilde{D})$ where $\tilde{D}<D\text{ and }\tilde{W}<W$. To ensure a fair comparison, we assume $\mathcal{O}(B\tilde{W}^{2}\tilde{D})\approx\mathcal{O}(W^{2}D)$ thereby keeping the representation capacity approximately the same across both architectures. However, for such separable architecture, the space complexity becomes exceedingly large as the complexity scales linearly with the dimension, $d$ of the PDE since $B=d$ leading to inefficiencies for high-dimensional problems. In comparison, Anant-Net also follows the separable architecture paradigm, with space complexity $\mathcal{O}(B\tilde{W}^{2}\tilde{D})$, but importantly with $B<<d$. As a result, the complexity of Anant-Net does not scale with dimensionality of the PDE, making it significantly more scalable.
We now analyze the time complexity of the proposed architecture. Let $N$ denote the number of collocation points sampled from a $d$-dimensional space for solving a $d$-dimensional PDE. Typically, $N$ is much larger than $d$. A standard fully connected neural network exhibits a time complexity of $\mathcal{O}(N^{d})$ for such problems. In contrast, separable architectures, such as SPINN [28], reduce the computational burden to $\mathcal{O}(Nd)$ by decomposing the multi-dimensional PDE into separable lower-dimensional components, thereby enabling more efficient training and inference. However, a key limitation of separable formulations becomes evident in extremely high-dimensional settings, where the linear growth in complexity with respect to the dimension $d$ remains prohibitive. The Anant-Net architecture addressed this challenge by tacking the high-dimensions batch-wise, achieving a time complexity of $\mathcal{O}\left(NB\right)$, where $B\ll d$ is a batch size. Therefore, the choice of $B$ emerges as an important hyperparameter that governs the trade-off between computational efficiency and model expressivity. Selecting an small value of $B$ (smaller than $d$) allows Anant-Net to maintain favorable scalability properties, especially in high-dimensional regimes. Importantly, the selection of $B$ may also be influenced by hardware constraints and data availability, both of which play an important role in determining the architecture’s practical effectiveness.
2.3 Universal Approximation Theorem: Through the lens of Anant-Net
In this section we define the universal approximation capabilities of Anant-Net architecture solving high-dimensional PDEs. For clarity, the Anant-Net ansatz is formulated using Einstein summation convention rather than explicit tensor product notation ($\otimes$).
Theorem 2.1
(Anant-Net) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$ we can find a sufficiently large $r>0$ (which is the size of the embedding layer for all body networks) and neural networks $f_{j}$ and $g_{j}$ such that,
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{j=1}^{r}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon$ (4)
$f_{j}=f_{j}(\hat{\mathcal{B}}_{k}^{1},X\text{\textbackslash}\hat{\mathcal{B}_{k}^{1})}\hskip 28.45274pt\text{and}\hskip 28.45274ptg_{j}=g_{j}(\hat{\mathcal{B}}_{k}^{2},Y\text{\textbackslash}\hat{\mathcal{B}}_{k}^{2})$ (5)
where, ^={^$\hat{\mathcal{B}}=\{\hat{\mathcal{B}}_{1},\hat{\mathcal{B}}_{2},\dots\hat{\mathcal{B}}_{k}\}$ is the set of sets consisting of active dimensions where at any given step-$k$, ^$\hat{\mathcal{B}}_{k}=\{\hat{\mathcal{B}}_{k}^{1},\hat{\mathcal{B}}_{k}^{2}\}=\{x_{p},x_{q}\}$ is the set of active dimensions, $\{X$\^$\hat{\mathcal{B}}_{k}^{1}\}$ and $\{Y$\^$\hat{\mathcal{B}}_{k}^{2}\}$ is the set of inactive dimensions. Here, $X=\{x_{1},x_{2},\dots x_{p}\dots x_{d/2}\}$, $Y=\{x_{d/2+1},\dots x_{q}\dots x_{d}\}$.
The proof is given in APPENDIX A.
2.4 Effect of Preconditioning on Training Dynamics
In this section, we investigate the role of preconditioning in enhancing the training of the Anant-Net for high-dimensional PDEs. Ryck et al. [31] established that linearly transforming the model parameters is effectively equivalent to preconditioning the loss gradient via multiplication with a positive definite matrix. Their findings demonstrated the utility of such preconditioning in accelerating convergence for a range of low-dimensional PDEs (typically in one or two spatial dimensions). Building on this foundation, the present study extends the theoretical framework to high-dimensional settings and demonstrates the effectiveness of gradient preconditioning in improving training performance when employing the Anant-Net architecture. It is important to acknowledge, however, that these benefits are accompanied by an increase in computational cost due to the cost associated with constructing and applying the preconditioning matrix. Consider a linear Anant-Net architecture with 3 body networks with the ansatz bearing the following form given by equation (1),
$\hat{u}_{\theta}(X_{1},X_{2},X_{3})=f_{1}^{(\theta_{1})}(X_{1})\otimes f_{2}^{(\theta_{2})}(X_{2})\otimes f_{3}^{(\theta_{3})}(X_{3}),$
where $f^{(\theta_{k})}(X_{k})=\sum_{i=1}^{n_{k}}w_{2,k}^{i}(w_{1,k}X_{k}+b_{1})=\sum_{i=1}^{n_{k}}w_{2,k}^{i}\phi_{i}(X_{k}),\leavevmode\nobreak\ k=1,2,3.$ The parameter $w^{i}_{j,k}$ denotes the weight associated with the $i^{\text{th}}$ neuron in the $j^{\text{th}}$ layer of the $k^{\text{th}}$ body network. In the above formulation, the basis functions are implicitly represented by the hidden layers of the body networks within the Anant-Net architecture, while the corresponding coefficients are extracted from the final output layer. This representation is analogous to the framework proposed by Cyr et al. [32], wherein a generic deep neural network is interpreted as a linear combination of learned basis functions. We consider the continuous formulation of the loss function, which serves as the analog of the discrete loss representation given in Equation (3), corresponding to the same PDE defined in Equation (2).
[…]
where, $x\in X\subset\mathbb{R}^{d}$. In this formulation, the Anant-Net architecture is approximated by a linear model (without any nonlinear activation functions), thereby yielding the linear representation presented above. For clarity and brevity, we restrict our analysis to the Anant-Net comprising three body networks, each containing two hidden layers. The resulting linear model is trained using standard gradient descent optimization. In its simplest form, the gradient descent update rule is given by $\theta_{r+1}=\theta_{r}-\eta\nabla L(\theta_{r}),$ where $\theta_{r}$ denotes the set of parameters at the $r^{\text{th}}$ iteration, and $\eta$ is the learning rate.
Given the inherent non-convexity of the loss function, an important aspect that makes analytical investigation particularly challenging, De Ryck et al. [31] propose an alternative formulation to simplify the analysis of gradient descent dynamics. Specifically, by fixing the parameters at a given iteration $r$, the Taylor series expansion of the network ansatz about the initialization point $\theta_{0}$ takes the following form:
$\hat{u}(x;\theta_{r})=\hat{u}(x;\theta_{0})+\nabla_{\theta}\hat{u}(x;\theta_{0})^{\mathsf{T}}(\theta_{r}-\theta_{0})+\frac{1}{2}(\theta_{r}-\theta_{0})^{\mathsf{T}}H_{r}(\theta_{r}-\theta_{0}),$
where $H_{r}$ is the hessian computed at intermediate parameters between 00 and $r$. Here, for brevity a new notation $\phi_{i}(x)=\partial_{\theta_{i}}u(x;\theta_{0})$ is introduced with $\mathcal{L}\phi_{i}\in L^{2}(\Omega)$. The above formulation is substituted in Equation 6 that leads to the following simplification:
$\theta_{r+1}=\theta_{r}-\eta\nabla L(\theta_{r})=(I-\eta\mathbb{A})\theta_{r}+\eta(\mathbb{A}\theta_{0}+\mathbb{B})+\eta\epsilon_{r}.$ (7)
Here, $\epsilon_{r}\in\mathbb{R}^{n}$ encapsulates the higher-order terms neglected in the Taylor series expansion, thereby isolating the leading-order contributions retained in the resulting expression for the simplified gradient descent formulation.
[…]
$\mathbb{B}_{i}=\langle f-\mathcal{L}u_{\theta_{0}},\mathcal{L}\phi_{i}\rangle_{L^{2}{(\Omega)}}+\lambda\langle u-u_{\theta_{0}},\phi_{i}\rangle_{L^{2}{(\Omega)}}\hskip 28.45274pt\mathbb{B}\in\mathbb{R}^{n}.$
In the case where the matrix $\mathbb{A}$ is invertible, let $\kappa(\mathbb{A})$ denote its condition number. To achieve a specified error tolerance $\epsilon$, the number of gradient descent iterations required satisfies the relation $N(\epsilon)\propto 1/\ln(1-c/\kappa(\mathbb{A}))$ for some constant $10<c<$ 1. Consequently, as $\mathbb{A}$ becomes increasingly ill-conditioned, i.e., $\kappa(\mathbb{A})\to\infty$, the number of iterations $N(\epsilon)$ diverges, indicating a deterioration in convergence rate. To address this challenge in the training of physics-informed machine learning models, we employ an explicit preconditioning strategy for the gradients, as detailed below.
$\widehat{\theta}_{r+1}=\widehat{\theta}_{r}-\eta\mathbb{PP}^{\mathsf{T}}\nabla L(\widehat{\theta}_{r}),$ (8)
where $\widehat{\theta}_{r}:=\mathbb{P}\theta_{r}$ and $\mathbb{P}\mathbb{P}^{\mathsf{T}}$ is a symmetric positive definite matrix. It follows from the preceding discussion that such a linear transformation of the parameters is mathematically equivalent to preconditioning the gradient. In this work, we adopt a preconditioning strategy grounded in (quasi-)Newton optimization methods, wherein the gradients are modified using curvature information derived from the loss landscape. Specifically, (quasi-)Newton approaches leverage both the gradient and an approximation of the Hessian of the loss function to accelerate convergence during optimization.
$\widehat{\theta}_{r+1}=\widehat{\theta}_{r}-\eta H^{-1}\nabla L(\widehat{\theta}_{r})$ (9)
[…]
$=\int_{\Omega}(\mathcal{L}\partial_{\theta_{i}}u_{\theta}(x))\cdot\mathcal{L}\partial_{\theta_{j}}u_{\theta}(x)dx+\lambda\int_{\partial\Omega}\partial_{\theta_{i}}u_{\theta}(x)\cdot\partial_{\theta_{j}}u_{\theta}(x)dx$
$=\int_{\Omega}(\mathcal{L}\phi_{i}(x)\cdot\mathcal{L}\phi_{j}(x)dx+\lambda\int_{\partial\Omega}\phi_{i}(x)\cdot\phi_{j}(x)dx=\langle\mathcal{L}\phi_{i},\mathcal{L}\phi_{j}\rangle_{L^{2}{(\Omega)}}+\lambda\langle\phi_{i},\phi_{j}\rangle_{L^{2}{(\Omega)}}$
$=\mathbb{A}_{i,j}$
By comparing Equation (8) and Equation (9) it can be concluded that using a (quasi-)Newton method implicitly preconditions the gradient since Hessian of the loss function is equivalent to $\mathbb{A}$.
Figure 3: Effect of Preconditioning on High-Dimensional Poisson Equation Training: Training results using (quasi-)Newton optimization for 6D, 9D, 15D, 18D, and 21D show overlapping performance without preconditioning. For 45D and higher (right), preconditioning offers no improvement, due to the linear architecture. Note that these experiments were performed on a T4 GPU with 15 GB RAM.
While preconditioning has demonstrated clear advantages in accelerating convergence within the framework of physics-informed deep learning, prior investigations have been primarily constrained to low-dimensional settings (typically $\leq$ 3 dimensions). In this work, we extend these findings to higher-dimensional PDEs by leveraging the proposed Anant-Net architecture. In particular, we employ a linear version of Anant-Net to validate theoretical predictions through empirical experiments that align closely with the theoretical analysis presented in this section. The corresponding results are illustrated in Figure 3. For clarity, we define high-dimensional as dimensionalities greater than 3 and up to 21, and extremely high-dimensional (X-high) as dimensionalities exceeding 21, as referenced in the figure. Within the high-dimensional regime, we observe that the theoretical benefits of preconditioning are preserved under the linear Anant-Net architecture. Specifically, preconditioning the gradients via Newton-based optimization yields stable training dynamics and favorable convergence behavior; see Figure 3 (left) . In contrast, omitting preconditioning causes the model to converge to suboptimal local minima, with negligible performance gains throughout training. However, in the X-high dimensional regime, the linear Anant-Net architecture fails to show any tangible benefit from preconditioning, as evidenced in Figure 3 (right). This breakdown of theoretical alignment in extremely high dimensions can be attributed to the inherent limitations of the linear two-layer network, which appears insufficient for capturing the complexity of the solution landscape in such settings. Furthermore, increasing the network’s depth or width does not overcome this limitation in the absence of nonlinearities. Nonetheless, as we elaborate in the Results section, these limitations can be effectively addressed by incorporating nonlinear activation functions, thereby enabling the resolution of high-dimensional PDEs. From a practical standpoint, to avoid the computational burden typically associated with second-order optimizers, we implement gradient preconditioning via (quasi-)Newton methods as a secondary optimization step. This step follows an initial training phase using a standard stochastic gradient-based optimizer.
2.5 Anant-Net with Interpretable KAN
Kolmogorov–Arnold neural networks (KANs), recently introduced by Liu et al. [33], extend the classical Kolmogorov–Arnold representation theorem, which was originally formulated for shallow networks with limited hidden neurons, to modern deep learning architectures. By integrating backpropagation-based training, the KAN framework generalizes naturally to networks with an arbitrary number of layers. In this formulation, a KAN with $L$ layers is expressed as a composition of $L$ basis functions (see Fig. 4):
$f(\mathbf{x})=(\Phi_{L-1}\circ\Phi_{L-2}\circ\dots\circ\Phi_{1}\circ\Phi_{0})(\mathbf{x}),$
where $f(\mathbf{x})$ is a multivariate continuous function with $\mathbf{x}\in\mathbb{R}^{d}$ and functionality of basis function $\Phi$ in each layer can be defined as follows:
$\mathbf{x}_{l+1}=\Phi_{l}\mathbf{x}_{l},$
where $\mathbf{x}_{l}\in\mathbb{R}^{n_{l}}$ and $\mathbf{x}_{l+1}\in\mathbb{R}^{n_{l+1}}$ denote the input and output of the $l$-th layer, respectively, in a KAN architecture with an arbitrary number of $L$ layers, and $n_{l}$ represents number of neurons in the $l$-th layer. The transformation $\Phi_{l}$ associated with the $l$-th layer can then be expressed in matrix form as follows:
$\Phi_{l}=\begin{bmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\dots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\dots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\dots&\phi_{l,n_{l+1},n_{l}}(\cdot)\\ \end{bmatrix}$
In both the original formulation [33] of the Kolmogorov–Arnold representation theorem and its recent extension ([34], [35], [36]), KANs were constructed using spline basis functions equipped with internal parameterizations. Two key hyperparameters govern the spline basis in this context: the polynomial order $k$ and the grid size $G$. These parameters play a central role in the performance of KANs. Notably, Theorem 2.1 in Liu et al. [33] establishes that the approximation error for any target function depends solely on $G$ and $k$, and remains independent of the input dimension $d$. More recently, several alternative basis functions have been proposed to enhance KANs, particularly for applications involving the solution of PDEs. In this work, we examine some of the most effective basis functions through a series of experiments. To the best of our knowledge, this study presents the first application of KANs to the solution of high-dimensional PDEs. Therefore, the efficiency of KANs in function approximation is fundamentally determined by the choice of basis functions $\Phi$. In this work, we investigate a range of basis formulations for KANs and systematically evaluate their performance in approximating the solution spaces of high-dimensional PDEs, using the Anant-Net architecture as a representative framework. KANs are particularly useful for their inherent interpretability, an attribute typically absent in standard multilayer perceptrons (MLPs). They are capable of learning both compositional structures and univariate functional mappings, often resulting in superior performance compared to MLPs. Furthermore, KANs have been shown to exhibit enhanced parametric efficiency, achieving comparable or improved predictive accuracy with significantly fewer trainable parameters than their MLP counterparts. Despite these advantages, several practical challenges remain unresolved. In particular, the selection of critical hyperparameters, such as grid size and spline order, is problem-dependent and often non-trivial, necessitating further investigation to enable more robust and automated deployments of KAN architectures.
Figure 4: Schematic of 3-layered Kolmogorov Arnold Networks (KAN) with input, $\mathbf{x}_{0}=\{x_{1},x_{2}\}$ and output, $f(\mathbf{x})=(\Phi_{2}\circ\Phi_{1}\circ\Phi_{0})(\mathbf{x}_{0})$. Replacing the ANN (body networks) in the Anant-Net shown in Figure 2 by plugging KAN (body networks) results in Anant-KAN architecture.
Recently, KANs have been employed for solving PDEs. Shukla et al. [36] introduced the PIKAN and DeepOKAN architectures, while Wang et al. [37] proposed Kolmogorov–Arnold Informed Neural Networks, which replace conventional MLP-based frameworks with KAN-based representations. Their study demonstrated that KANs, when coupled with lower-order orthogonal polynomials instead of the originally proposed spline-based basis functions, exhibit superior performance across a range of problems. Despite this success, some models using KAN with orthogonal polynomials experienced instability during training, though the cause of these divergences remains unclear. Furthermore, their work was primarily limited to low-dimensional PDEs. In a similar vein, PIKAN was extended to a separable framework, leading to the development of SPIKAN by Jacob et al. [38]. The advantages of this separable framework were demonstrated on specific problems, such as the 2D Helmholtz equation, with the maximum dimensionality considered in their work being $\leq 3$. It is important to note that KAN has the potential to address the curse of dimensionality (see Theorem 2.1 in Liu et al. [33]), which asserts that the approximation error of KAN for a $d$-dimensional function $f$ is independent of $d$. We aim to investigate the immediate benefits of this property by extending Anant-Net to solve high-dimensional PDEs. To the best of our knowledge, the advantages of KAN for high-dimensional PDEs have not yet been fully explored, and we seek to investigate these benefits through the Anant-KAN (Anant-Net with KAN) architecture.
2.5.1 Computational Complexity for Anant-KAN
In this context, we consider space complexity as a function of the number of trainable parameters in a given model architecture for KANs. For a KAN with depth $D$ (number of layers) and width $W$ (number of nodes per layer), the space complexity of the model can be approximated as $\mathcal{O}(W^{2}D(G+k))$, which simplifies to $\mathcal{O}(W^{2}DG)$ for a spline basis with grid size $G$ and spline order $k$. However, for a separable KAN (such as SPIKAN [38]) with $B$ body networks, the space complexity is given by $\mathcal{O}(B\tilde{W}^{2}\tilde{D}G)$, where $\tilde{D}<D$ and $\tilde{W}<W$. To ensure a fair comparison, we assume that $\mathcal{O}(B\tilde{W}^{2}\tilde{D}G)\approx\mathcal{O}(W^{2}DG)$, thereby maintaining approximately the same representation capacity across both architectures. However, in such a separable architecture, the space complexity scales linearly with the dimensionality $d$ of the PDE, since $B=d$, leading to inefficiencies for high-dimensional problems. In contrast, Anant-KAN also follows the separable architecture paradigm, with space complexity $\mathcal{O}(B\tilde{W}^{2}\tilde{D}G)$ using a spline basis with grid size $G$, but crucially with $B\ll d$. As a result, the complexity of Anant-KAN does not scale with the dimensionality $d$ of the PDE, making it significantly more scalable. It is important to note that the computational time complexity of Anant-KAN is similar to that of Anant-Net, which is $\mathcal{O}(NB)$, with $B\ll d$ for solving high-dimensional PDEs.
3 Results
This section presents the empirical investigations performed using Anant-Net to solve high-dimensional linear and nonlinear PDEs over hypercubes with domain length $L>1$, where the volume increases with the number of dimensions. Specifically, we focus on the Poisson equation, sine-Gordon equation, and Allen-Cahn equation. A key challenge in such settings is the enforcement of boundary conditions, where boundary sampling significantly contributes to the complexity of the problem as the dimensionality increases. To evaluate the performance of the proposed architecture, we use the percentage relative $L_{2}$ error, defined as:
[…]
Uncertainty quantification is performed by reporting the test error over 10 independent random realizations, with data points sampled from the high-dimensional domain. Unless otherwise specified, all experiments are executed on a T4 GPU with 15 GB RAM. To this end, we systematically compare the performance of Anant-Net with other state-of-the-art methods that also employ boundary sampling strategies, rather than enforcing boundary conditions directly. The weights for the individual loss terms is fixed for all the cases: $\lambda_{r}=1$ and $\lambda_{b}=15$ in Equation 3.
In addition to the architectural hyperparameters in deep neural networks, several algorithmic hyperparameters must be tuned to achieve optimal performance with the Anant-Net architecture. These include: the number of body networks ($B$), grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#$\#CG$), number of boundary/data grids (#$\#BG$), and the sampling frequency ($S_{\text{FREQ}}$). The number of body networks ($B$) directly influences the architecture of Anant-Net. For example, when $B=3$, the architecture consists of three body networks that process 3D cubes sampled from the high-dimensional space. Similarly, when $B=4$, four body networks are used to process 4D tesseracts from the high-dimensional domain. The sampling frequency ($S_{\text{FREQ}}$) directly affects the training process by determining how frequently a new set of dimensions is sampled during training. If $S_{\text{FREQ}}$ is too low, the model spends less time interacting with the newly sampled subspace, leading to poor generalization. Conversely, if $S_{\text{FREQ}}$ is too high, the model may overfit due to prolonged exposure to the same subspace. Finally, the grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#$\#CG$), and number of boundary/data grids (#$\#BG$) determine the size and structure of the dataset handled by Anant-Net. We will employed $B=3$ for all computational experiments unless stated otherwise.
Figure 5: Point-wise absolute testing error on randomly selected 3D cubes from 21D, 99D, and 300D cases for the (left) Poisson equation, (center) sine-Gordon equation, and (right) Allen–Cahn equation. Here, ”$\mathsf{ActiveDims}$: $\{p,q,r\}$” denotes the active dimensions in the $d$-dimensional space where gradients are non-zero, and $\mathrm{D}$\$\{p,q,r\}$ represents the inactive dimensions. $\mathrm{D}=\{1,2,\dots,d\}$ denotes the set of all dimensions.
3.1 High-dimensional Poisson’s equation
The high-dimensional Poisson’s equation is a fundamental PDE that arises in various scientific and engineering disciplines, particularly in modeling steady-state processes such as electrostatics, heat conduction, and gravitational potentials in complex systems. In high-dimensional settings, such as those encountered in quantum mechanics and financial mathematics, the computational complexity of solving the equation increases exponentially. Here, we consider Poisson’s equation of the following form:
$-\Delta u(x)=f(x),x\in\Omega,$ (10)
$u(x)=g(x),x\in\partial\Omega,$ (11)
where $\Omega\in[-1,1]^{d}$ and the exact solution is assumed to be of the form:
$u_{\text{exact}}(x)=\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}^{2}+\sin\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)},$
and
$f(x)=\frac{1}{d}\biggl{(}\sin\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}-2\biggl{)}.$
Dimension Architecture Optimizer ##\##Collocation ##\##Data
21D $\bigl{[}7,64,64,10\bigr{]}\times 3$ Adam+L-BFGS 38416 6912
99D $\bigl{[}33,64,64,40\bigr{]}\times 3$ Adam+L-BFGS 90552 10800
300D $\bigl{[}100,128,128,128\bigr{]}\times 3$ Adam+L-BFGS 274400 32400
Table 2: Anant-Net architecture used for solving high-dimensional Poisson equation
Figure 6: Poisson Equation: (Left) Training time profiling; (Right) mean testing error with standard deviation, computed over 10 random realizations (seeds), for the 21D, 99D, and 300D cases.
In this section, we solve the high-dimensional Poisson’s equation in 21D, 99D, and 300D. For all cases, we use the Adam optimizer to train Anant-Net, followed by L-BFGS to further fine-tune the learning process. The benefits of preconditioning the gradients were discussed in previous sections, where we empirically validated the theory using (quasi-)Newton optimization with a linear version of Anant-Net. However, in this section, Anant-Net incorporates a non-linear activation function ($\tanh$) to solve the high-dimensional Poisson’s equation. Table 2 provides further details on the specific hyperparameters used. As suggested by the curse of dimensionality, the number of collocation and data samples required to solve high-dimensional PDEs on a hypercube grows exponentially with dimension. This poses a major challenge for existing methods, including state-of-the-art PINNs such as NTK-based PINNs [29], self-adaptive PINNs [39], etc., which struggle with scalability due to the computational burden of large sample volumes. Anant-Net addresses this issue efficiently. Table 2 summarizes the total number of collocation and boundary points used for solving Poisson’s equation in 21D, 99D, and 300D. The sample sizes are controlled via hyperparameters: $N_{C}$, #$\#CG$ for collocation points, and $N_{B}$, #$\#BG$ for boundary points. Figure 6 (left) indicates that boundary data sampling is the primary contributor to training time across all dimensions, highlighting a need to optimize this component. Meanwhile, Figure 6 (right) shows a mean relative $L_{2}$ testing error along with the standard deviation, evaluated on randomly selected test points from the high-dimensional space. Figure 5 (left) shows the absolute point-wise error on randomly selected 3D cubes from the 21D, 99D, and 300D cases. More details on the hyperparameters used for training is explained in APPENDIX B.
3.2 High-dimensional sine-Gordon equation
The high-dimensional sine-Gordon equation is a nonlinear PDE that generalizes the classical sine-Gordon model to multiple spatial dimensions. It serves as a canonical model in mathematical physics, with relevance to field theory, nonlinear wave propagation, and condensed matter systems. In higher dimensions, the equation admits complex solution structures, including multidimensional solitons and breathers, whose computation is impeded by both the nonlinearity and the high-dimensional nature of the problem. We consider sine-Gordon equation of the following form:
$-\Delta u(x)+\sin(u(x))=f(x),x\in\Omega,$
$u(x)=g(x),x\in\partial\Omega,$
where $\Omega\in[-1,1]^{d}$ and the exact solution is assumed to be of the form:
$u_{\text{exact}}(x)=\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}^{2}+\sin\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)},$
and
$f(x)=\frac{1}{d}\biggl{(}\sin\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}-2\biggl{)}+\sin(u_{\text{exact}}(x)).$
Dimension Architecture Optimizer ##\##Collocation ##\##Data
21D $\bigl{[}7,64,64,10\bigr{]}\times 3$ Adam+L-BFGS 54880 12960
99D $\bigl{[}33,64,64,40\bigr{]}\times 3$ Adam+L-BFGS 109760 14040
300D $\bigl{[}100,64,64,128\bigr{]}\times 3$ Adam+L-BFGS 301840 34560
Table 3: Anant-Net architecture used for solving high-dimensional sine-Gordon equation
Figure 7: Sine-Gordon Equation: (Left) Training time profiling; (Right) mean testing error with standard deviation, computed over 10 random realizations (seeds), for the 21D, 99D, and 300D cases.
As opposed to the previous section, where Anant-Net was applied to a linear PDE, we now solve a nonlinear PDE; the high-dimensional sine-Gordon equation in 21D, 99D, and 300D. Except for the 300D case, the architectural and optimization hyperparameters remain consistent with the previous setup. Due to the nonlinearity of the sine-Gordon equation, the number of collocation and data samples is slightly higher compared to the Poisson equation. Details of the data sampling are provided in Table 3. This increase is justified, as the model must account for both the growing hypercube volume with dimension and the added complexity introduced by nonlinearity. Additionally, for the sine-Gordon equation, we employ a layer-wise adaptive activation function of the form proposed in [40]. Figure 7 (left) shows that boundary data sampling remains the primary contributor to training time, consistent with the findings from the previous section. Figure 7 (right) shows a mean relative $L_{2}$ testing error along with the standard deviation, evaluated on randomly selected test points from the high-dimensional space. The testing error increases with the dimensionality. Figure 5 (middle) shows the absolute point-wise error on randomly selected 3D cubes from the 21D, 99D, and 300D cases. More details on the hyperparameters used for training is explained in APPENDIX B.
3.3 High-dimensional Allen-Cahn equation
The high-dimensional Allen–Cahn equation is a nonlinear reaction-diffusion PDE that models phase separation phenomena in multi-component systems. Initially formulated to describe the motion of anti-phase boundaries in binary alloys, its extension to high-dimensional domains has garnered significant interest for capturing complex interfacial dynamics and pattern formation in physical, chemical, and biological systems. We consider Allen-Cahn equation of the following form:
$-\Delta u(x)+u(x)-u(x)^{3}=f(x),x\in\Omega,$
$u(x)=g(x),x\in\partial\Omega,$
where $\Omega\in[-1,1]^{d}$ and the exact solution is assumed to be of the form:
$u_{\text{exact}}(x)=\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}^{2}+\sin\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)},$
and
$f(x)=\frac{1}{d}\biggl{(}\sin\biggl{(}\frac{1}{d}\sum_{i=1}^{d}x_{i}\biggl{)}-2\biggl{)}+u_{\text{exact}}(x)-u_{\text{exact}}(x)^{3}.$
Dimension Architecture Optimizer ##\##Collocation ##\##Data
21D $\bigl{[}7,64,64,10\bigr{]}\times 3$ Adam+L-BFGS 38416 8640
99D $\bigl{[}33,64,64,40\bigr{]}\times 3$ Adam+L-BFGS 264000 36450
300D $\bigl{[}100,64,64,128\bigr{]}\times 3$ Adam+L-BFGS 274400 43200
Table 4: Anant-Net architecture used for solving high-dimensional Allen-Cahn equation
Figure 8: Allen-Cahn Equation: (Left) Training time profiling; (Right) mean testing error with standard deviation, computed over 10 random realizations (seeds), for the 21D, 99D, and 300D cases.
The high-dimensional Allen–Cahn equation is solved in 21D, 99D, and 300D with minimal modifications to the architecture and optimization hyperparameters used in the previous sections. Details on hyperparameters and data sampling are provided in Table 4. Although we again address a nonlinear PDE, in this case we employ a fixed nonlinear activation function ($\sin$) without adaptive parameters. As illustrated in Figure 8 (left), the performance trends align with earlier observations: boundary data sampling remains the dominant contributor to the overall computational cost during training. An exception is observed in the same figure, where the training times for the 21D and 99D cases are notably lower compared to similar problems discussed in previous sections. This reduction in runtime is likely attributable to the fewer training iterations required to achieve comparable accuracy for the Allen–Cahn equation when using Anant-Net. Figure 8 (right) shows the mean and standard derivation of relative $L_{2}$ testing error for 21D, 99D, and 300D cases. Figure 5 (right) shows the absolute point-wise error on randomly selected 3D cubes from the 21D, 99D, and 300D cases. Additional details on the number of training iterations used across all high-dimensional PDEs are provided in the APPENDIX B.
3.4 Effect of Boundary Data Sampling on Anant-Net Accuracy
The optimality of boundary data sampling reflects a fundamental trade-off between model accuracy and computational efficiency. While a large volume of boundary data can improve prediction accuracy, it often incurs substantial computational costs during training. Conversely, reducing the volume of boundary data lowers computational demand but may compromise model accuracy. Striking a balance between these two factors is particularly important when solving high-dimensional PDEs. In this section, we examine the sensitivity of Anant-Net’s accuracy to varying volumes of boundary data in high-dimensional PDE problems. Although test accuracy may also depend on factors such as network architecture, the quantity and quality of boundary data are expected to have a significant influence. Results are presented for both linear and nonlinear non-homogeneous high-dimensional PDEs using clean data.
[…]
In particular, we consider the 300-dimensional Poisson and sine-Gordon equations. In this study, the volume of boundary data is varied while keeping the volume of collocation points, network architecture, and all other hyperparameters fixed for Anant-Net. Profiling of the training time (Fig. 9) confirms that variations in boundary data volume are the primary contributors to changes in computational cost. As shown in Fig. 9, for both linear and nonlinear high-dimensional PDEs, the test accuracy, evaluated on randomly sampled test points, initially improves with increased boundary data but tends to plateau beyond a certain threshold. For the 300-dimensional Poisson and sine-Gordon equations, training is performed using an initial sweep with the Adam optimizer followed by L-BFGS. For the linear high-dimensional Poisson equation, a clear trend is observed wherein increased boundary data enhances model accuracy, albeit at the cost of longer training times (Fig. 9). A similar pattern is evident for the nonlinear high-dimensional sine-Gordon equation. These results indicate that, regardless of the underlying PDE type, increasing the boundary data volume improves the predictive accuracy of Anant-Net, but at the expense of higher computational cost.
3.5 Anant-KAN Results
Recent studies have explored various basis functions for function approximation using KANs, targeting both continuous functions and PDE solutions. Table 5 highlights efficient bases for PDE solution spaces. We extend these bases to high-dimensional PDEs using our proposed architecture, Anant-Net. Theorem 2.1 in [33] shows that KAN approximation error depends only on grid size $G$ and spline order $k$, not on input dimension, suggesting its ability to overcome the curse of dimensionality; validated empirically in this work. While [33] emphasizes splines as universal bases via the Kolmogorov-Arnold theorem, alternative bases like Chebyshev polynomials have also shown promise. Chebyshev-KAN [35] benefits from orthogonality and recursive computation, improving training stability and efficiency. However, instability was noted for low-dimensional PDEs in [36], causing training errors to diverge. This issue was addressed by introducing Modified Chebyshev-KAN, which incorporates $\tanh$ activations in all but the final layer to suppress high-frequency instabilities.
Network Method Basis $(\phi)$ Activation
Anant-KAN1 KAN [33] Spline -
Anant-KAN2 Chebyshev KAN [35] Chebyshev Polynomials Tanh(1)
Anant-KAN3 Modified Chebyshev KAN [36] Chebyshev Polynomials Tanh(2)
Anant-KAN4 Fourier KAN [34] Fourier ($a.cos(kx)+b.sin(kx)$) -
Table 5: Variants of KAN with different basis functions $(\phi)$. (1) represents tanh activation applied only at the input layer for normalization while (2) represents tanh activation applied to all layer except the last layer. Anant-KAN for different basis functions for $\bigl{[}7,64,64,10\bigr{]}\times 3$ configuration of architecture.
Figure 10: Comparison of Anant-KAN with different basis functions for solving 21D Poisson equation and sine-Gordon equation for a fixed order of basis $(k)=3$.
In this section, we focus on enhancing the interpretability of our proposed architecture by leveraging the KAN, referred to here as Anant-KAN. This is the first application of KAN to high-dimensional PDEs. One of the primary advantages of KAN lies in its parameter efficiency, as demonstrated in Figure 10. However, despite its compact parameterization, KAN exhibits significant computational overhead, which persists regardless of the dimensionality of the PDE. Our computational experiments further corroborate this limitation, especially in high-dimensional settings.
Figure 11: Sensitivity of polynomial order ($k$) and grid size ($G$) on relative $L_{2}$ error and run-time for 21D Poisson equation using Anant-KAN with spline basis (Anant-KAN1).
The key findings of the Anant-KAN architecture are as follows:
  • For both linear and non-linear PDEs, spline basis functions consistently demonstrate superior approximation capability in high-dimensional PDEs using Anant-KAN framework. This performance trend is clearly illustrated in Fig. 10.
  • Building on this observation, we further refine our investigation by identifying the optimal configuration of the spline basis, characterized by the pair $(k,G)$, denotes the polynomial order and the grid resolution, respectively. The sensitivity of Anant-KAN to variations in these parameters is presented in Fig. 11.
  • In general, for a fixed value of grid size $G$, increasing the polynomial order $k$ improves the accuracy of the Anant-KAN. Also, accuracy becomes more sensitive to changes in $k$ at larger values of $G$.
  • In contrast, runtime analysis reveals that increasing $k$ for any fixed $G$ incurs a substantial computational cost, as shown in the Fig. 11 (right). Consequently, for high-dimensional PDEs, a parameter configuration of $k\leq 3$ and $G\leq 10$ makes an effective balance between accuracy and computational efficiency.
  • We apply the Anant-KAN architecture to solve high-dimensional PDEs using the configuration $k=2$ and $G=5$, coupled with additional hyperparameter tuning strategies such as adaptive activation functions [41, 40, 42, 43]. The resulting improvements in predictive performance are shown in Fig. 12.
  • Finally, we observe a distinct increase in relative $L_{2}$ error for the high-dimensional Poisson equation as the grid resolution $G$ increases. This phenomenon is likely attributable to overfitting, a hypothesis supported by earlier observations in [33], where an abrupt rise in testing error occurred beyond a certain grid size, termed the interpolation threshold. While this behavior was originally reported in low-dimensional problems, our study demonstrates its early onset in high-dimensional PDE contexts, suggesting the need for regularization or adaptive complexity control in such settings.
The high computational cost associated with KAN imposes a significant constraint on training strategies for Anant-KAN and similar architectures. In particular, the use of second-order optimization methods such as L-BFGS becomes computationally prohibitive for high-dimensional PDEs, necessitating reliance on first-order optimizers such as Adam. This limitation is clearly reflected in our computational experiments. Among the various basis functions evaluated, the spline basis consistently provides the most accurate representation of solutions to high-dimensional PDEs, as shown in Fig. 10. On the other hand, the Fourier basis (Anant-KAN4) performs the poorest across both linear and non-linear PDEs.
Figure 12: Comparison of Anant-KAN with spline basis (order ($k$) = 2, grid size ($G$) = 5) for 21D and 99D Poisson equation, sine-Gordon equation, and Allen-Cahn equation using layer-wise adaptive-sin activation [40] for all body networks. Note that the mean relative $L_{2}$ error on random samples with standard deviation was obtained from 10 random realizations (seeds).
The Anant-KAN architecture, which employes spline basis functions, is further enhanced through fine-tuning, enabling accurate solutions of both linear and nonlinear high-dimensional partial differential equations, as demonstrated in Figure 12. These results strongly suggest that the proposed Anant-KAN architecture can achieve exceptional accuracy for a broad class of PDEs. The point-wise absolute error associated with Anant-KAN, evaluated over a randomly sampled set of dimensions from the high-dimensional space, is presented in Figure 13. All experiments in this section are conducted on a high-RAM CPU system with a 50 GB RAM configuration. Nonetheless, the implementation can be further optimized for GPU acceleration to enable parallelism and scale Anant-KAN to even higher-dimensional problems.
Figure 13: Point-wise absolute testing error for random 3D hyper-surfaces of 21D, 99D (left) Poisson equation (center) sine-Gordon equation, (right) Allen-Cahn equation. Here, ”$\mathsf{ActiveDims}$: $(p,q,r)$” represents the active dimensions in the $d$-dimensional space where gradients are non-zero and $\mathrm{D}-\{p,q,r\}$ are inactive. Also, $\mathrm{D}=\{1,2,\dots,d\}$ is a set which constitutes all the dimensions.
3.6 Comparison
In this section, we compare the testing performance and runtime of various methods. Specifically, we focus on methods that are designed to solve high-dimensional PDEs in hypercube domains with $L>1$ employing boundary data sampling strategy. We report both the testing error and the runtime on GPU and CPU platforms.
Method Device Run time (min) (%) Relative $L_{2}$ Error
21D 99D 300D 21D 99D 300D
WAN [26] CPU 250 - - 0.9 $\pm$ 0.08 NaN NaN
Anant-Net CPU 14 53 366 4.5 $\pm$ 0.037 8.6 $\pm$ 0.038 4.5 $\pm$ 0.037
PINN GPU# 34 100 250 16.5 $\pm$ 0.099 100.6 $\pm$ 0.90 174.5 $\pm$ 0.15
Anant-Net GPU# 25 69 127 1.5 $\pm$ 0.013 7.1 $\pm$ 0.037 6.07 $\pm$ 0.038
Table 6: Run time and (%) relative $L_{2}$ testing error on randomly selected test points for high-dimensional Poisson equation. (High RAM CPU with 51 GB RAM; #T4 GPU with 15 GB RAM). The best testing error is shown in bold numbers.
Method Device Run time (min) (%) Relative $L_{2}$ Error
21D 99D 300D 21D 99D 300D
WAN [26] CPU 167 - - 1.21 $\pm$ 0.05 NaN NaN
Anant-Net CPU 41 100 147 5.8 $\pm$ 0.030 7.4 $\pm$ 0.053 7.9 $\pm$ 0.025
PINN GPU# 37 124 297 49.1 $\pm$ 0.32 35.3 $\pm$ 0.23 246.6 $\pm$ 0.35
Anant-Net GPU# 53 108 168 2.7 $\pm$ 0.024 6.3 $\pm$ 0.052 6.7 $\pm$ 0.041
Table 7: Run time and (%) relative $L_{2}$ error on randomly selected test points for high-dimensional sine-Gordon equation. (High RAM CPU with 51 GB RAM; #T4 GPU with 15 GB RAM). The best testing error is shown in bold numbers.
As shown in Table 6 and Table 7, WAN [26] demonstrates the efficiency of their method for problems with dimensions less than 50. For moderately high dimensions (e.g., 21D), WAN achieves better accuracy compared to Anant-Net, while Anant-Net attains significantly higher computational efficiency for the same volume of training data. Although WAN maintains stability and efficiency at lower dimensions, it still suffers from the curse of dimensionality at higher dimensions. For example, WAN becomes unstable at larger dimensions (99D and 300D), with the error diverging to NaN (not a number) during training. This instability is consistently observed regardless of the optimizer type, network architecture, or volume of training data. Furthermore, Table 6 and Table 7 indicate that Anant-Net is both stable and computationally efficient at larger dimensions, whereas WAN is unstable and incurs high computational cost. The studies presented in Table 6 and Table 7 were conducted to ensure a fair comparison, although the methods and architectures differ in various respects. For example, the number of parameters was kept approximately the same across all models despite differences in their architectures. The weak form based on WAN [26] was trained using the AdaGrad optimizer for both the primal and adversarial networks, as proposed in their paper. In contrast, the Anant-Net and PINN was trained using the Adam optimizer, with the pretrained model subsequently fine-tuned using the L-BFGS optimizer with appropriate learning rate selection. L-BFGS is a quasi-Newton optimizer that approximates the Hessian to avoid the high computational cost of direct computation for large problems. However, in non-convex settings, poor Hessian initialization can cause a mismatch between the true and approximated Hessians, leading to performance drops or optimizer divergence. Rafati et al. [44] showed that batch size and Hessian initialization significantly influence L-BFGS performance, even though their study focused on the MNIST dataset. We believe these insights extend to solving high-dimensional PDEs. Thus, investigating batch size and initialization strategies for L-BFGS in the context of Anant-Net is a promising direction for future work.
4 Conclusions
In this work, we introduced Anant-Net, a novel and scalable neural architecture tailored for solving high-dimensional partial differential equations (PDEs) defined over hypercubic domains. Traditional numerical and learning-based methods often struggle with the curse of dimensionality, especially when the side length of the domain $L>1$, resulting in significant computational bottlenecks. Anant-Net overcomes these limitations through several key innovations, summarized below:
  • Anant-Net employs a tensor product formulation that enables efficient computation across batch dimensions. This design allows for dimension-wise sweeps over the full batch of collocation points, enabling the model to solve PDEs in hundreds of dimensions with high accuracy.
  • To address the computational cost of automatic differentiation in high-dimensional spaces, Anant-Net adopts a selective differentiation approach. Instead of exhaustively computing derivatives in all dimensions, it samples a subset of dimensions (equal to batch-size) and applies differentiation only where needed. This significantly reduces computational overhead while preserving the essential structure of the governing PDEs.
  • The model demonstrates the ability to solve both linear and nonlinear PDEs in up to 300 dimensions within a few hours on a single GPU. We report the mean and standard deviation (variance) of relative $L_{2}$ testing error on randomly selected test points from the high-dimensional spaces. Furthermore, the accuracy can be enhanced by increasing the number of training and collocation points.
  • We also propose an interpretable variant of Anant-Net, based on the Kolmogorov–Arnold Network (KAN) for high-dimensional PDEs. This variant offers deeper insights into the learned solution structure and enhances interpretability in deep neural networks.
For extremely high-dimensional problems where the data volume exceeds single-GPU memory, Anant-Net supports efficient training through mini-batching. Moreover, it can be easily adapted to parallelization techniques such as data and model parallelism [45], tensor parallelism [14], and domain decomposition [46, 47, 48], enabling deployment across multi-GPU architectures. Anant-Net presents a promising architecture for solving high-dimensional problems with both high accuracy and computational efficiency.
Appendix A Universal Approximation Theorem: Through the lens of Anant-Net
Theorem A.1
(Anant-Net) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$, we can find a sufficiently large $r>0$ and neural networks $f_{j}$ and $g_{j}$ such that,
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{j=1}^{r}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon$ (12)
$f_{j}=f_{j}(\hat{\mathcal{B}}_{k}^{1},X\text{\textbackslash}\hat{\mathcal{B}}_{k}^{1})\hskip 28.45274pt\text{and}\hskip 28.45274ptg_{j}=g_{j}(\hat{\mathcal{B}}_{k}^{2},Y\text{\textbackslash}\hat{\mathcal{B}}_{k}^{2})$ (13)
where, ^={^$\hat{\mathcal{B}}=\{\hat{\mathcal{B}}_{1},\hat{\mathcal{B}}_{2},\dots\hat{\mathcal{B}}_{k}\}$ is the set of sets consisting of active dimensions where at any given step-$k$, ^$\hat{\mathcal{B}}_{k}=\{\hat{\mathcal{B}}_{k}^{1},\hat{\mathcal{B}}_{k}^{2}\}=\{x_{p},x_{q}\}$ is the set of active dimensions, $\{X\text{\textbackslash}\hat{\mathcal{B}}_{k}^{1}\}$ and $\{Y\text{\textbackslash}\hat{\mathcal{B}}_{k}^{2}\}$ is the set of inactive dimensions. Here, $X=\{x_{1},x_{2},\dots x_{p}\dots x_{d/2}\}$, $Y=\{x_{d/2+1},\dots x_{q}\dots x_{d}\}$.
Proof. Let ${\phi_{i}}$ and ${\psi_{j}}$ be the orthonormal basis for $L^{2}(X)$ and $L^{2}(Y)$ respectively. Then ${\phi_{i}\psi_{j}}$ forms the orthonormal basis for $L^{2}(X\times Y)$. Therefore, we can find a sufficiently large r, consider an active dimension set $\mathcal{B}_{k}$ that results in the solution $u_{k}$ such that,
$\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}\biggr{\|}_{L^{2}(X\times Y)}<\frac{\varepsilon}{2}$ (14)
where $a^{k}_{ij}$ denotes
[…]
On the other hand, by the universal approximation theorem, we can find neural networks $f_{i}$ and $g_{j}$ such that,
$\|\phi_{i}-f_{j}\|_{L^{2}(X)}\leq\frac{\varepsilon}{3^{j}\|u^{k}\|_{L^{2}(X\times Y)}}\text{ and }\|\psi_{j}-g_{j}\|_{L^{2}(Y)}\leq\frac{\varepsilon}{3^{j}\|u^{k}\|_{L^{2}(X\times Y)}}$ (16)
We consider the difference between $u^{k}$ and $\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}$ by adding and subtracting $\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}$ as shown below:
$\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}\biggr{\|}_{L^{2}(X\times Y)}=\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}+\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}-\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}\biggr{\|}_{L^{2}(X\times Y)}$
Appropriately grouping the terms and applying the triangular inequality $\|p+q\|\leq\|p\|+\|q\|$, we get the following.
$\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}\biggr{\|}_{L^{2}(X\times Y)}\leq\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}\biggr{\|}_{L^{2}(X\times Y)}+\biggl{\|}\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}\biggr{\|}_{L^{2}(X\times Y)}$
But from equation, equation can be further simplified to the following form.
$\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}\biggr{\|}_{L^{2}(X\times Y)}\leq\frac{\varepsilon}{2}+\biggl{\|}\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}\biggr{\|}_{L^{2}(X\times Y)}$
Now we go ahead to quantify the below unknown to complete the proof.
$\biggl{\|}\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}\biggr{\|}_{L^{2}(X\times Y)}$
Consider the terms inside the parenthesis as follows,
$\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}\psi_{j}-\sum_{i,j}^{r}a^{k}_{ij}f_{i}g_{j}=\sum_{i,j}^{r}a^{k}_{ij}\phi_{i}(\psi_{j}-g_{j})-\sum_{i,j}^{r}a^{k}_{ij}(\phi_{i}-f_{i})g_{j}=II_{1}+II_{2}=II$
We first compute $II_{1}$.
[…]
We apply Cauchy-Scharwz inequality, $|\langle p,q\rangle|\leq\|p\|\cdot\|q\|$ to further break-down the above expression as follows.
$\|II_{1}\|^{2}_{L^{2}(X\times Y)}\leq\biggl{(}\int_{X}\sum_{j}^{r}|A_{j}(x)|^{2}dx\biggr{)}\biggl{(}\int_{Y}\sum_{j}^{r}|B_{j}(y)|^{2}dy\biggr{)}$
Then,
$\int_{X}\sum_{j}^{r}|A_{j}(x)|^{2}dx=\int_{X}\sum_{j}^{r}\biggl{(}\sum_{i}^{r}a^{k}_{i,j}\phi_{i}\biggr{)}^{2}dx$
$\hskip 62.59596pt=\sum_{j}^{r}\biggl{\|}\sum_{i}^{r}a^{k}_{i,j}\phi_{i}\biggr{\|}^{2}_{L(X)}$
$\hskip 56.9055pt=\sum_{j}^{r}\biggl{(}\sum_{i}^{r}|a^{k}_{i,j}|\biggr{)}^{2}$
$\hskip 56.9055pt\leq 2\sum_{j}^{r}\sum_{i}^{r}|a^{k}_{i,j}|^{2}<2\sum_{i,j}^{\infty}|a^{k}_{i,j}|^{2}=2\|u^{k}\|^{2}_{L^{2}(X\times Y)}$
On the other hand, we have the following form.
[…]
$\hskip 56.9055pt=\sum_{j}^{r}\|\psi_{j}(y)-g_{j}(y)\|^{2}_{L^{2}(Y)}$
$\hskip 56.9055pt\leq\sum_{j}^{r}\frac{\varepsilon^{2}}{9^{j}\|u^{k}\|^{2}_{L^{2}(X\times Y)}}$
$\hskip 56.9055pt<\frac{\varepsilon^{2}}{8\|u^{k}\|^{2}_{L^{2}(X\times Y)}}$
Hence we have.
$\|II_{1}\|^{2}_{L^{2}(X\times Y)}<\frac{\varepsilon^{2}}{8\|u^{k}\|^{2}_{L^{2}(X\times Y)}}2\|u^{k}\|^{2}_{L^{2}{(X\times Y)}}=\frac{\varepsilon^{2}}{4}$
Likewise,
$\|II_{2}\|^{2}_{L^{2}(X\times Y)}<\frac{\varepsilon^{2}}{4}$
Therefore,
$\|II\|^{2}_{L^{2}(X\times Y)}<\frac{\varepsilon^{2}}{2}$
We go back to with estimates,
$\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon$
However, we know that the approximation error is an average of the error across all the sets of active dimensions and accounting for $k$ sets of active dimensions results in the following.
$\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\sum_{k}^{|\hat{\mathcal{B}}|}\varepsilon=|\hat{\mathcal{B}}|\cdot\varepsilon$
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i,j}^{r}a^{k}_{i,j}f_{i}g_{j}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
Since, the summation of the product of $f_{i}$ and $g_{j}$ is performed element-wise, it is sensible to reorder the notations as shown below.
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}b_{i}\tilde{f_{i}}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
where, $b_{i}$ is a constant and grouping it with $\tilde{f_{i}}$ gives the following.
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}\{b_{i}\tilde{f_{i}}\}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}=\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{\tilde{r}}t_{i}\tilde{g_{i}}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
Replacing, $\tilde{r}$, $\tilde{g_{i}}$, $t_{i}$ with $r$, $g_{i}$, $f_{i}$ results in the following form.
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{i}^{r}f_{i}g_{i}\biggr{\|}_{L_{2}(X\times Y)}<\varepsilon$
Hence proved $\blacksquare$
Appendix B Hyperparameters for Anant-Net
The main hyperparameters for Anant-Net include: the number of body networks ($B$), grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#$\#CG$), number of boundary/data grids (#$\#BG$), and the sampling frequency ($S_{\text{FREQ}}$). The hyperparameters used for the main results presented in this paper are discussed in this section. For clarity, the tables presented below follow the format: ##\##Collocation $=(N_{C}\times N_{C}\times N_{C})\times\#CG$ and ##\##Data $=(N_{B}\times N_{B}\times N_{B})\times\#BG$ for the current cases dealing with 3 body networks. However, if any future work deals with 4 body networks, then ##\##Collocation $=(N_{C}\times N_{C}\times N_{C}\times N_{C})\times\#CG$ and ##\##Data $=(N_{B}\times N_{B}\times N_{B}\times N_{B})\times\#BG$ and likewise.
Hyperparameters for Section 3.1
Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
21D $\bigl{[}7,64,64,10\bigr{]}\times 3$ Adam 5000 70000
L-BFGS 5000 70000
99D $\bigl{[}33,64,64,40\bigr{]}\times 3$ Adam 5000 165000
L-BFGS 5000 165000
300D $\bigl{[}100,128,128,128\bigr{]}\times 3$ Adam 500 50000
L-BFGS 2000 200000
Table 8: Anant-Net architecture used for solving high-dimensional Poisson equation
Dimension ##\##Collocation ##\##Data
21D $(14\times 14\times 14)\times 14=38416$ $(6\times 6\times 6)\times 32=6912$
99D $(14\times 14\times 14)\times 33=90552$ $(6\times 6\times 6)\times 50=10800$
300D $(14\times 14\times 14)\times 100=274400$ $(6\times 6\times 6)\times 150=32400$
Table 9: Details on collocation and data points sampled for cases in Table 8 for solving high-dimensional Poisson equation
Hyperparameters for Section 3.2
Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
21D $\bigl{[}7,64,64,10\bigr{]}\times 3$ Adam 5000 100000
L-BFGS 5000 100000
99D $\bigl{[}33,64,64,40\bigr{]}\times 3$ Adam 5000 200000
L-BFGS 5000 200000
300D $\bigl{[}100,64,64,128\bigr{]}\times 3$ Adam 1000 110000
L-BFGS 2000 140000
Table 10: Anant-Net architecture used for solving high-dimensional sine-Gordon equation
Dimension ##\##Collocation ##\##Data
21D $(14\times 14\times 14)\times 20=54880$ $(6\times 6\times 6)\times 60=12960$
99D $(14\times 14\times 14)\times 40=109760$ $(6\times 6\times 6)\times 65=14040$
300D $(14\times 14\times 14)\times 110=301840$ $(6\times 6\times 6)\times 160=34560$
Table 11: Details on collocation and data points sampled for cases in Table 10 for solving high-dimensional sine-Gordon equation
Hyperparameters for Section 3.3
Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
21D $\bigl{[}7,64,64,10\bigr{]}\times 3$ Adam 5000 70000
L-BFGS 5000 70000
99D $\bigl{[}33,64,64,40\bigr{]}\times 3$ Adam 3000 99000
L-BFGS 5000 165000
300D $\bigl{[}100,64,64,128\bigr{]}\times 3$ Adam 500 50000
L-BFGS 1000 200000
Table 12: Anant-Net architecture used for solving high-dimensional Allen-Cahn equation
Dimension ##\##Collocation ##\##Data
21D $(14\times 14\times 14)\times 14=38416$ $(6\times 6\times 6)\times 40=8640$
99D $(20\times 20\times 20)\times 33=264000$ $(9\times 9\times 9)\times 50=36450$
300D $(14\times 14\times 14)\times 100=274400$ $(6\times 6\times 6)\times 200=43200$
Table 13: Details on collocation and data points sampled for cases in Table 12 for solving high-dimensional Allen-Cahn equation
Hyperparameters for Section 3.5
Equation Dimension Architecture Optimizer Sampling Frequency ##\##Iterations
Poisson 21D $\bigl{[}7,5,5,10\bigr{]}\times 3$ Adam 5000 70000
99D $\bigl{[}33,5,5,14\bigr{]}\times 3$ Adam 2000 80000
sine-Gordon 21D $\bigl{[}7,5,5,10\bigr{]}\times 3$ Adam 5000 70000
99D $\bigl{[}33,5,5,14\bigr{]}\times 3$ Adam 2000 80000
Allen-Cahn 21D $\bigl{[}7,5,5,10\bigr{]}\times 3$ Adam 5000 70000
99D $\bigl{[}33,5,5,14\bigr{]}\times 3$ Adam 2000 80000
Table 14: Anant-KAN architecture used for solving high-dimensional Poisson equation, sine-Gordon equation and Allen-Cahn equation. Note that Spline basis is used for all the cases with polynomial order ($k$) = 2 and grid size ($G$) = 5
Equation Dimension ##\##Collocation ##\##Data
Poisson 21D $(14\times 14\times 14)\times 14=38416$ $(6\times 6\times 6)\times 32=6912$
99D $(14\times 14\times 14)\times 40=109760$ $(6\times 6\times 6)\times 65=14040$
sine-Gordon 21D $(14\times 14\times 14)\times 14=38416$ $(6\times 6\times 6)\times 32=6912$
99D $(14\times 14\times 14)\times 40=109760$ $(6\times 6\times 6)\times 50=10800$
Allen-Cahn 21D $(14\times 14\times 14)\times 14=38416$ $(6\times 6\times 6)\times 32=6912$
99D $(14\times 14\times 14)\times 40=109760$ $(6\times 6\times 6)\times 65=14040$
Table 15: Details on collocation and data points sampled for cases in Table 14 for solving high-dimensional Poisson equation, sine-Gordon equation and Allen-Cahn equation.
Appendix C Nomenclature
Symbol Definition
$\mathcal{L}$ Differential operator for PDE
$\mathcal{B}$ Boundary operator
$f$ Source term of the PDE
$g$ Boundary source term of the PDE
$u$ Exact solution of the PDE
$\hat{u}_{\theta}$ Neural Network (NN) ansatz
$\theta_{i}$ Trainable NN parameters
$\mathrm{D}$ Dimension set $\{1,2,\dots d\}$ for a $d-$dimensional PDE
$\lambda_{i}$ Weight on loss component-$i$
$x_{i}$ Input along dimension-$i$
$X_{i}$ Input set of dimensional features $\{x_{1},x_{2}\dots x_{d/3}\}$ for body network-$i$
$\{p,q,r\}$ A stochastically sampled set of active dimensions where p, q, r are assigned to body network 1, 2, 3 respectively
$\mathcal{G}$ Grid of boundary/collocation
$\mathcal{J}$ Loss function
$N$ Number of collocation points
$B$ Number of body networks
$r$ Size of the embedding layer in each body network
$H$ Hessian of the loss function
$\Omega$, $\partial\Omega$ Domain, Domain Boundary
$f_{i}$, $g_{j}$ Body Networks in Anant-Net as mentioned in Section 2.3 on UAT
$\hat{\mathcal{B}}$ Set of active dimension sets as mentioned in Section 2.3 on UAT
Table 16: Definition of symbols used in the paper